# RL3: Boosting Meta Reinforcement Learning via RL inside RL2

### v4

@@ 4.1 Theoretical Justification
Proof: Given a task distribution $\mathcal{M}$, then for state $s$, there exists a maximum object-level optimal value function $V^{*}_{max}(s)$, corresponding to some MDP $M_{max}\in\mathcal{M}$, such that for all MDPs $M_{i}\in\mathcal{M}$, $V^{*}_{max}(s)\geq V^{*}_{i}(s)$. Observe that the expected cumulative discounted reward experienced by the agent cannot be greater than the most optimistic value function over all tasks, since $\bar{V}^{*}(\bar{b})$ is a weighted average of individual value functions $V^{\pi_{\theta}}(s)$, which are themselves upper bounded by $V^{*}_{max}(s)$. Thus,
@@
Next, we see that combining the asymptotic accuracy of Q-estimates and Equation (2) gives us
@@
Equation (4) (see full proof in Appendix A.3) shows that for $t\geq\kappa$, acting greedily w.r.t. $Q_{i}^{*}$ leads to Bayes-optimal behavior, and knowing the Bayes-optimal belief over tasks is not required, which also implies that the entire experience history can be ignored at that point. Moreover, it follows from equation (4) that for $t<\kappa$,
@@
where error $\varepsilon_{i}(\Upsilon)$ is the error in Q-value estimates, which is a function of the dynamics of MDP $M_{i}$ and data (both amount and quality) seen by the agent so far in the experience history $\Upsilon$. While this error will diminish as $t\rightarrow\infty$, in the short run, a function $f(\Upsilon)$ could be learned to either estimate the error or estimate $\bar{V}^{*}(\bar{b})$ entirely.
@@
Finally, we note that near-perfect function approximation of $\bar{V}^{*}(\bar{b})$ as $t\rightarrow\infty$ reduces error in meta-value function approximation for all preceding belief states, as meta-values for consecutive belief states $\bar{b}$ and $\bar{b}^{\prime}$ are linked through the Bellman equation for BAMDPs (see details in Appendix A.3):
This dependency helps in meta-training in RL3, as most (outer) RL algorithms used for training the meta-critic have a temporal-difference based learning component. Without conditioning on Q-estimates, error in $\bar{V}^{*}(\bar{b})$ would instead increase as $t\rightarrow\infty$, as the meta-critic would be conditioned on a larger history, which could destabilize the meta-value learning for all preceding belief states during meta-training.
Algorithm 1 Value-Augmenting Wrapper for Discrete MDPs
procedure ResetMDP(vamdp)
    vamdp.$t\leftarrow 0$; vamdp.$t_{\tau}\leftarrow 0$
    vamdp.rl $\leftarrow\textsc{InitRL}()$
    $s=\textsc{ResetMDP}$(vamdp.mdp)
    return $\textsc{OneHot}(s)\cdot Q[s]\cdot N[s]$
procedure StepMDP(vamdp, $a$)
    $s\leftarrow$ mdp.$s$
    $r,s^{\prime}\leftarrow\textsc{StepMDP}$(vamdp.mdp, $a$)
    $d\leftarrow\textsc{Terminated}$(vamdp.mdp)
[…]
    vamdp.$Q\leftarrow\textsc{UpdateRL}$(vamdp.rl, $s,a,r,s^{\prime},d$)
[…]
         vamdp.$t_{\tau}\leftarrow 0$
         $s^{\prime}\leftarrow\textsc{ResetMDP}$(vamdp.mdp)
    return $r,$ $\textsc{OneHot}(s^{\prime})\cdot Q[s^{\prime}]\cdot N[s^{\prime}]$
procedure Terminated(vamdp)
    return vamdp.$t\geq H$
@@ 4.2 Implementation
Implementing RL3 involves simply replacing each MDP in the task distribution with a corresponding value-augmented MDP (VAMDP) and solving the resulting VAMDP distribution using RL2. Each VAMDP has the same action space and reward function as the corresponding MDP. The value augmented state $\hat{s}_{t}\in S\times\mathbb{R}^{k}\times\mathbb{I}^{k}$ includes the object level state $s_{t}$, $k$ real values and $k$ integer values for the Q-estimates ($Q^{t}(s_{t},a)$) and action counts ($N^{t}(s_{t},a)$) for each of the $k$ actions. When the object-level state space $S$ is discrete, $s_{t}$ needs to be represented as an $|S|$-dimensional one-hot vector. Note that the value augmented state space is continuous. In the VAMDP transition function, the object-level state $s$ has the same transition dynamics as the original MDP, while the dynamics of Q-estimates are a function of $T$, $R$, and the specific object-level RL algorithm used for estimating Q-values. An episode of the VAMDP spans the entire interaction period with the corresponding MDP, which may include multiple episodes of the MDP, as Q-estimates continue to evolve beyond episode boundaries. In code, a VAMDP RL environment is implemented as a wrapper over a given MDP environment. The pseudocode is provided in Algorithm 1 and additional implementation details and hyperparameters for RL2 and RL3 are mentioned in Appendix B.
@@ 5 Experiments
We compare RL3 to a modified version of RL2. First, we replace LSTMs with transformers in both the meta-actor and meta-critic for the purpose of mapping experiences to actions and meta-values, respectively. This is done to improve RL2 ’s ability to handle long-term dependencies instead of suffering from vanishing gradients. Moreover, RL2 -transformer trains significantly faster than RL2 -LSTM. Second, we include in the state space the total number of interaction steps and the total number of steps within each episode during a meta-episode (see Fig. 1). Third, we use PPO Schulman et al. (2017) for training the meta actor-critic, instead of TRPO Schulman et al. (2015). These modifications and other minor-implementation details incorporate the recommendations made by Ni et al. (2022), who show that model-free recurrent RL is competitive with other state-of-the-art meta RL approaches such as VeriBAD Zintgraf et al. (2020), if implemented properly. RL3 simply applies the modified version of RL2 to the distribution of value-augmented MDPs explained in section 4.2. Within each VAMDP, our choice of object-level RL is a model-based algorithm to maximize data efficiency – we estimate a tabular model of the environment and run finite-horizon value-iteration using the model.
@@
In our test domains, each meta-episode involves procedurally generating an MDP according to a parameterized distribution, which the meta-actor interacts with for a fixed interaction budget $H$. This interaction might consist of multiple object-level episodes of variable length, each of which are no longer than a maximum task horizon. For a given experiment, each approach is trained on the same series of MDPs. For testing, each approach is evaluated on an identical set of 1000 MDPs distinct from the training MDPs. For testing OOD generalization, MDPs are generated from distributions with different parameters than in training. We select three task domains for our experiments, which cover a range of short-term, long-term, and complex dependencies.
@@
Bernoulli Bandits: We use the same setup described in Duan et al. (2016) with $k=5$ arms. To test OOD generalization, we generate bandit tasks by sampling success probabilities from $\mathcal{N}(0.5,0.5)$.
@@
Random MDPs: We use the same setup described in Duan et al. (2016). The MDPs have 10 states and 5 actions with stochastic transitions and rewards, and the task horizon is 10. The mean rewards and transition probabilities are drawn from a normal and a flat Dirichlet distribution, respectively. To test OOD generalization, the rewards are generated deterministically and initialized from $\mathcal{U}(0,2)$.
GridWorld Navigation: A set of navigation tasks in a 2D grid environment. We experiment with 11x11 (121 states) and 13x13 (169 states) grids. The agent always starts in the center of the grid and needs to navigate through obstacles to a single goal location. The grid also contains slippery tiles, fatally dangerous tiles and warning tiles surrounding the latter. See Fig. 3(a) for an example of a 13x13 grid. Instead of one-hot vectors, we use the 2D $(x,y)$ grid location to represent agent state. To test OOD generalization, we vary parameters including the stochasticity of actions, density of obstacles and the number of dangerous tiles. For this domain, we consider an additional variation of RL3, called RL3 -coarse where a given grid is partitioned into clusters of tiles (or abstract states), each of size 2. Abstract states are comprised only of adjacent, traversable cells, and are used solely for the purpose of estimating the object-level Q-values. Our goal is to test whether coarse-level Q-value estimates are still useful to the meta-RL policy. The domains and the abstraction strategy used for the RL3 -coarse approach are described in greater detail in Appendices E and B.3, respectively.
6 Results
In this section, we present several experiments that demonstrate the effectiveness of RL3 . Beyond matching or exceeding the performance of RL2 in all test domains, RL3 also shows better OOD generalization, which we attribute to the increased generality of the Q-value representation. More striking, the advantages of RL3 increase significantly with longer interactions periods and less stochastic tasks. We hypothesize that this is due to the increased accuracy of the Q-value estimates in these cases. Last, we find that RL3 performs well even with coarse-grained object-level RL over abstract states, showing minimal drop in performance in most cases and even occasional increases, while the computational savings are substantial.
We emphasize that the core of our approach, which is augmenting MDP states with action-value estimates, is not inherently tied to RL2 and is orthogonal to most other meta-RL research. VAMDPs can be plugged into any base meta-RL algorithm with a reasonable expectation of improving it.
Bandits: Table 2 shows the results for the Bandits domains. For $H=100$ and $H=500$, both approaches perform comparably. However, the OOD generalization for RL3 is slightly better. We also experiment with a Markovian version of RL3, where we use a feed-forward neural network that is conditioned only on the Q-estimates and action-counts, since those are sufficient for Bayes-optimal behavior in this domain. As expected, the results are similar to those with regular RL3 .
MDPs: Table 3 shows the results for the MDPs domains. Once again, for $H=100$ and $H=500$, both approaches perform comparably, and once again, OOD generalization is slightly better for RL3 . We suspect that for such short budgets on this highly stochastic domain, Q-estimates do not converge enough to be very useful. To test this hypothesis, we test both approaches by applying the models trained for $H=500$ to $H=1000$ (row 4 in Table 3) using a moving window context for the transformer models (see Appendix B for details). Here, RL3 generalizes significantly better, demonstrating the utility of Q-estimates when they are allowed to converge for more iterations. In fact, the score achieved by RL3 when trained for $H=500$ and tested on $H=1000$ is similar to that of the original RL2 implementation Duan et al. (2016) when trained specifically for $H=1000$.
GridWorlds: Table 4 shows the results for the GridWorld Navigation domain. On 11x11 grids with $H=250$, RL3 significantly outperforms RL2. On 13x13 grids with $H=350$, the outperformance margin is even greater, showing that while RL2 struggles with a greater number of states, a longer adaptation period and more long-term dependencies, RL3 can take advantage of the Q-estimates to overcome the challenge. We also test the OOD generalization of both approaches in different ways by increasing the obstacle density (Dense), making actions on non-water tiles deterministic (Deterministic), increasing the number of wet ‘W’ tiles (Watery), increasing the number of danger ‘X’ tiles (Dangerous) and having the goal only in the corners (Corner). On all variations, RL3 continues to significantly outperform RL2. In a particularly interesting outcome, both approaches show improved performance on the Deterministic variation. However, RL3 gains 80% more points than RL2 does, which is likely because Q-estimates converge faster on this less stochastic MDP and therefore provide greater help to RL3 . Conversely, in the Watery variation, which is more stochastic, both RL2 and RL3 lose roughly equal number of points. Finally, in each case, RL3 -coarse significantly outperforms RL2. In fact, it performs on par with RL3, even outperforming it on Corner variation, except on the canonical 13x13 case and its Deterministic variation, where it scores about 90% of the scores for RL3 .
@@ 6 Results
Table 3: Test scores (mean $\pm$ standard error) for MDPs domain and the OOD variation. +Time limit doubled without retraining policy.
Budget $H$ RL2 RL3
100 $159.5\pm 0.8$ $158.9\pm 0.8$
500 $927.8\pm 3.7$ $926.9\pm 3.7$
500 $772.8\pm 1.7$ $\mathbf{775.9}\pm 1.7$
1000+ $1871.8\pm 7.4$ $\mathbf{1916.8}\pm 7.4$
Table 4: Test scores (mean $\pm$ standard error) for GridWorlds domain and the OOD variations (13x13, $H$=350).
Variation RL2 RL3 RL3 -coarse
11x11, $H$=250 $524\pm 22$ $\mathbf{630}\pm 22$ $\mathbf{613}\pm 21$
13x13, $H$=350 $584\pm 28$ $\mathbf{902}\pm 27$ $831\pm 28$
Dense $384\pm 26$ $\mathbf{691}\pm 28$ $\mathbf{674}\pm 28$
Deterministic $960\pm 38$ $\mathbf{1574}\pm 35$ $1464\pm 36$
Watery $513\pm 26$ $\mathbf{826}\pm 27$ $\mathbf{822}\pm 28$
Dangerous $283\pm 28$ $\mathbf{647}\pm 30$ $\mathbf{657}\pm 30$
Corner $319\pm 23$ $508\pm 23$ $\mathbf{645}\pm 23$
@@
Fig. 3 shows a sequence of snapshots of a meta-episode where the trained RL3 agent is interacting with an instance of a 13x13 grid. The first snapshot shows the agent just before reaching the goal for the first time. Prior to the first snapshot, the agent had explored many locations in the grid. The second snapshot shows the next episode just after the agent finds the goal, resulting in value estimates being updated using object-level RL for all visited states. Snapshot 3 shows the agent consequently using the Q-estimates to navigate to the goal presumably by choosing high-value actions. The agent also explores several new nearby states for which it does not have Q-estimates. Snapshot 4 shows the final Q-value estimates. A short video of the GridWorld environment, showing both RL2 and RL3 agents solving a set of problem instances side-by-side, is included in the supplementary material.
@@
Computation Overhead Considerations: As mentioned earlier, for implementing object-level RL, we use model estimation followed by finite-horizon value-iteration to obtain Q-estimates. The computation overhead is negligible for Bandits (5 actions, task horizon = 1) and very little for the MDPs domain (10 states, 5 actions, task horizon 10). For 13x13 GridWorlds (up to 169 states, 5 actions, task horizon = 350), RL3 takes approximately twice the computation time of RL2 per meta-episode. However, RL3 -coarse requires only 10% overhead while still outperforming RL2 and retaining more than 90% of the performance of RL3 . This demonstrates the utility of state abstractions in RL3 for scaling. Our implementation is available in the supplementary material.
7 Conclusion

### v5

@@ 4.1 Theoretical Justification
Proof: Given a task distribution $\mathcal{M}$, then for state $s$, there exists a maximum object-level optimal value function $V^{*}_{max}(s)$, corresponding to some MDP $M_{max}\in\mathcal{M}$, such that for all MDPs $M_{i}\in\mathcal{M}$, $V^{*}_{max}(s)\geq V^{*}_{i}(s)$. The expected cumulative discounted reward experienced by the agent cannot be greater than the most optimistic value function over all tasks, since $\bar{V}^{*}(\bar{b})$ is a weighted average of individual value functions $V^{\pi_{\theta}}(s)$, which are themselves upper bounded by $V^{*}_{max}(s)$. Thus,
@@
Next, we see that combining the asymptotic accuracy of Q-estimates and Equation (2) yields
@@
Equation 4 (proof in Appendix A.3) shows that for $t\geq\kappa$, acting greedily w.r.t. $Q_{i}^{*}$ leads to Bayes-optimal behavior, and knowing the Bayes-optimal belief over tasks is not required, implying that the experience history can be ignored at that point. Moreover, it follows from equation 4 that for $t<\kappa$,
@@
where error $\varepsilon_{i}(\Upsilon)$ is the error in Q-value estimates. While this error will diminish as $t\rightarrow\infty$, in the short run, a function $f(\Upsilon)$ could be learned to either estimate the error or estimate $\bar{V}^{*}(\bar{b})$ entirely.
@@
Finally, we note that near-perfect function approximation of $\bar{V}^{*}(\bar{b})$ as $t\rightarrow\infty$ reduces error in meta-value function approximation for all preceding belief states, as meta-values for consecutive belief states $\bar{b}$ and $\bar{b}^{\prime}$ are linked through the Bellman equation for BAMDPs (see details in Appendix A.3)
$\bar{V}^{*}(\bar{b})=\max_{a\in A}\Big{[}\!\sum_{M_{i}\in\mathcal{M}}\!\!\bar{b}(i)R_{i}(s,a)+\gamma\!\sum_{\bar{\omega}\in\bar{\Omega}}\bar{O}(\bar{\omega}|\bar{b},a)\bar{V}^{*}(\bar{b}^{\prime})\Big{]}.$ (6)
This dependency helps meta-training in RL3 with temporal-difference based learning algorithms. Without conditioning on Q-estimates, error in $\bar{V}^{*}(\bar{b})$ would instead increase as $t\rightarrow\infty$, as the meta-critic would be conditioned on a larger history, which could destabilize the meta-value learning for all preceding belief states during meta-training.
@@ 4.2 Implementation
Implementing RL3 involves simply replacing each MDP in the task distribution with a corresponding value-augmented MDP (VAMDP) and solving the resulting VAMDP distribution using RL2. Each VAMDP has the same action space and reward function as the corresponding MDP. The value augmented state $\hat{s}_{t}\in S\times\mathbb{R}^{k}\times\mathbb{I}^{k}$ includes the object level state $s_{t}$, $k$ real values and $k$ integer values for the Q-estimates ($Q^{t}(s_{t},a)$) and action counts ($N^{t}(s_{t},a)$) for each of the $k$ actions.In practice, we provide action advantages along with the max Q-value (value function) instead of Q-estimates. When the object-level state space $S$ is discrete, $s_{t}$ needs to be represented as an $|S|$-dimensional one-hot vector. Note that the value augmented state space is continuous. In the VAMDP transition function, the object-level state $s$ has the same transition dynamics as the original MDP, while the dynamics of Q-estimates are a function of $T$, $R$, and the specific object-level RL algorithm used for estimating Q-values. An episode of the VAMDP spans the entire interaction period with the corresponding MDP, which may include multiple episodes of the MDP, as Q-estimates continue to evolve beyond episode boundaries. In code, a VAMDP RL environment is implemented as a wrapper over a given MDP environment. The pseudocode, additional implementation details and hyperparameters for RL2 and RL3 are mentioned in Appendix
@@ 5 Experiments
We compare RL3 to our enhanced implementation of RL2. In our implementation, we replace LSTMs with transformers in both the meta-actor and meta-critic for the purpose of mapping experiences to actions and meta-values, respectively. This is done to improve RL2 ’s ability to handle long-term dependencies instead of suffering from vanishing gradients. Moreover, RL2-transformer trains significantly faster than RL2-LSTM. Second, we include in the state space the total number of interaction steps and the total number of steps within each episode during a meta-episode (see Fig. 1). Third, we use PPO [2] for training the meta actor-critic, instead of TRPO [54]. These modifications and other minor-implementation details incorporate the recommendations made by Ni et al. [6], who show that model-free recurrent RL is competitive with other state-of-the-art meta RL approaches such as VariBAD [7], if implemented properly. RL3 simply applies the modified version of RL2 to the distribution of value-augmented MDPs explained in section 4.2. Within each VAMDP, our choice of object-level RL is a model-based algorithm to maximize data efficiency – we estimate a tabular model of the environment and run finite-horizon value-iteration using the model. Once again, we emphasize that the core of our approach, which is augmenting MDP states with action-value estimates, is not inherently tied to RL2 and is orthogonal to most other meta-RL research. VAMDPs can be plugged into any base meta-RL algorithm with a reasonable expectation of improving it.
@@
In our test domains, each meta-episode involves procedurally generating an MDP according to a parameterized distribution, which the meta-actor interacts with for a fixed adaptation period, or interaction budget, $H$. This interaction might consist of multiple object-level episodes of variable length, each of which are no longer than a maximum task horizon. For a given experiment, each approach is trained on the same series of MDPs. Each experiment is done for 3 seeds and the results of the median performing model are reported. For testing, each approach is evaluated on an identical set of 1000 MDPs distinct from the training MDPs. For testing OOD generalization, MDPs are generated from distributions with different parameters than in training. We select three discrete domains for our experiments, which cover a range of short-term, long-term, and complex dependencies. These domains both reflect the challenges faced by meta-RL and simultaneously allow transparent analysis of the results.
@@
Bernoulli Bandits: We use the same setup described by Duan et al. [4] with $k=5$ arms. To test OOD generalization, we generate bandit tasks by sampling success probabilities from $\mathcal{N}(0.5,0.5)$. We should note that this is an easy domain and serves as a sanity check to ensure that Q-value estimates do not hurt RL3, causing inferior performance.
@@
Random MDPs: We use the same setup described by Duan et al. [4]. The MDPs have 10 states, 5 actions, and task horizon 10. The rewards and transition probabilities are drawn from a normal and a flat Dirichlet distribution ($\alpha=1.0$), respectively. OOD test MDPs use Dirichlet $\alpha=0.25$. We should note that this domain is particularly challenging for RL3 due to the high degree of stochasticity and thus the slower convergence rate of Q-estimates.
@@
GridWorld Navigation: A set of navigation tasks in a 2D grid environment. We experiment with 11x11 (121 states) and 13x13 (169 states) grids. The agent starts in the center and needs to navigate through obstacles to a single goal. The grid also contains slippery tiles, dangerous tiles and warning tiles. See Fig. 4(a) for an example of a 13x13 grid. The state representation is coordinates $(x,y)$. To test OOD generalization, we vary parameters including the stochasticity of actions, density of obstacles and the number of dangerous tiles. For this domain, we consider an additional variation of RL3, called RL3-coarse where a given grid is partitioned into clusters of 2 adjacent tiles (or abstract states), which are used solely for the purpose of estimating the object-level Q-values. Our goal is to test whether coarse-level Q-value estimates are still useful to the meta-RL policy. The domains and the abstraction strategy are described in greater detail in Appendices E and B.3, respectively.
6 Results
In summary, we observe that beyond matching or exceeding the performance of RL2 in all test domains i) RL3 shows better OOD generalization, which we attribute to the increased generality of the Q-value representation, ii) the advantages of RL3 increase with longer interactions periods and less stochastic tasks, which we attribute to the increased accuracy of the Q-value estimates, iii) RL3 performs well even with coarse-grained object-level RL over abstract states with substantial computational savings, showing minimal drop in performance in most cases, and iv) RL3 shows faster meta-training.
Figure 3: Results for the MDPs and GridWorlds domains. Figure 3a shows the average cumulative reward (negligible standard error) earned as a fraction of the oracle policy for in-distribution (solid) and OOD (dashed) tasks; Figure 3b shows the fraction of RL2-transformer meta-training iterations that RL3 requires (variance is insignificant across seeds) to match RL2-transformer performance or fully converge, both as functions of the adaptation period. Note the log horizontal axis on both plots. Figure 3c shows the average cumulative reward ($\pm$ standard error) earned by RL2, RL3 , and RL3-coarse agents on several variations of the GridWorlds domain.
Bandits: Fig 2 shows the results for this sanity-check domain. For $H=100$ and $H=500$, both approaches perform comparably. However, the OOD generalization for RL3 is slightly better. We also experiment with a Markovian version of RL3, where a feed-forward neural network is conditioned only on the Q-estimates and action-counts, since those are sufficient for Bayes-optimal behavior in this domain. As expected, the results are similar to regular RL3 .
MDPs: Figures 3a and 3b show the results for the MDPs domain. In Figure 3a, we see that for relatively short budgets, $H\leq 500$, both RL3 and RL2-transformer perform comparably on in-distribution problems, with RL3 performing slightly better on OOD tasks. We suspect that, due to the short budgets and highly stochastic domain, Q-estimates do not converge enough to be very useful for RL3 . However, as the budget increases, we see that RL3 continues to improve while RL2-transformer actually becomes worse and the performance gap on both in-distribution and OOD tasks becomes significant. Overall, we see that RL3 preserves asymptotic scaling properties of traditional RL while simultaneously maintaining strong OOD performance. Moreover RL3 it is able to learn meta-policies much more efficiently. Figure 3b shows the number of iterations of PPO RL3 takes to converge completely, as well as to match the performance of RL2-transformer, measured as a fraction of the time it takes for RL2-transformer to converge. This advantage of RL3 is again most pronounced for longer adaptation periods, but we still do observe significant meta-training speedup on even moderate ones. Overall, it is clear that as adaptation periods grow, RL3 achieves nearer-to-optimal policies in a fraction of the meta-training time and maintains better OOD generalization.
GridWorlds: Fig 3c shows the results for the GridWorld domain. On 11x11 grids with $H=250$, RL3 significantly outperforms RL2. On 13x13 grids with $H=350$, the performance margin is even greater, showing that while RL2-transformer struggles with a greater number of states, a longer adaptation period and more long-term dependencies, RL3 can take advantage of the Q-estimates to overcome the challenge. We also test the OOD generalization of both approaches in different ways by varying certain parameters of the 13x13 grids, namely, increasing the obstacle density (Dense), making actions on non-water tiles deterministic (Deterministic), increasing the number of wet ‘W’ tiles (Watery), increasing the number of danger ‘X’ tiles (Dangerous) and having the goal only in the corners (Corner). On all variations, RL3 continues to significantly outperform RL2. In a particularly interesting outcome, both approaches show improved performance on the Deterministic variation. However, RL3 gains 80% more points than RL2 , which is likely because Q-estimates converge faster on this less stochastic MDP and therefore provide greater help to RL3 . Conversely, in the Watery variation, which is more stochastic, both RL2 and RL3 lose roughly equal number of points. Overall, in each case, RL3-coarse significantly outperforms RL2-transformer. In fact, it performs on par with RL3, even outperforming it on Corner variation, except on the canonical 13x13 case and its Deterministic variation, where it scores about 90% of the scores for RL3. Finally, we see similar meta-training speedups where RL3 requires just $50\%$ and $30\%$ of the total iterations to match the performance of RL2-transformer on the 11x11 and 13x13 grids, respectively.
@@ 6 Results
Fig. 4 shows a sequence of snapshots of a meta-episode where the trained RL3 agent is interacting with an instance of a 13x13 grid. The first snapshot shows the agent just before reaching the goal for the first time. Prior to the first snapshot, the agent had explored many locations in the grid. The second snapshot shows the next episode just after the agent finds the goal, resulting in value estimates being updated using object-level RL for all visited states. Snapshot 3 shows the agent consequently using the Q-estimates to navigate to the goal presumably by choosing high-value actions. The agent also explores several new nearby states for which it does not have Q-estimates. Snapshot 4 shows the final Q-value estimates. A set of short videos of the GridWorld environment, showing both RL2 and RL3 agents solving the same set of problem instances, can be found at https://youtu.be/eLA_S1BQUYM.
@@
Computation Overhead Considerations: As mentioned earlier, for implementing object-level RL, we use model estimation followed by finite-horizon value-iteration to obtain Q-estimates. The computation overhead is negligible for Bandits (5 actions, task horizon = 1) and very little for the MDPs domain (10 states, 5 actions, task horizon 10). For 13x13 GridWorlds (up to 169 states, 5 actions, task horizon = 350), RL3 takes approximately twice the computation time of RL2 per meta-episode. However, RL3-coarse requires only 10% overhead while still outperforming RL2 and retaining more than 90% of the performance of RL3. This demonstrates the utility of state abstractions in RL3 for scaling. Finally, the meta-training sample efficiency demonstrated by RL3 translates directly to wall-time efficiency as training is dominated by gradient computation, not value iteration during data collection in PPO. Our implementation is available at https://github.com/bhatiaabhinav/RL3.
7 Limitations and Conclusion
Though it compares favorably to strong meta-RL approaches like RL2-transformer where applicable, RL3 does have some limitations. First, it assumes the object-level decision-making model is an MDP, which although a common assumption in the literature, may be challenged in practice. While in principle we could extend RL3 to POMDPs using methods like point-based value iteration, this has yet to be tested empirically. Second, RL3 relies on fast, potentially approximate methods for object-level RL, and using value iteration complicates application to problems with continuous state spaces. However, we speculate that a crude linear function approximation would suffice. Finally, inference time is slightly slower at deploy time due to running object-level RL. However, the overall training time is actually faster because of better meta-training efficiency. In fact, RL3 could enable working with adaptation periods that are otherwise prohibitively long for many meta-RL approaches.